OpenStudy (anonymous):

will give fan & medal :) Choose the correct description of the graph of the compound inequality: x − 3 < −7 or x + 5 greater than or equal to 8 A number line with an open circle on −4, shading to the left, and a closed circle on 3, shading to the right A number line with a closed circle on −4, shading to the left, and an open circle on 3, shading to the right A number line with an open circle on −4, a closed circle on 3, and shading in between A number line with a closed circle on −4, an open circle on 3, and shading in between

OpenStudy (decentnabeel):

first split it up x-3<-7 and x+5>=8 x-3<-7 +3 +3 x<-4 for the second part x+5>=8 -5 -5 x>=3 so it would end up being x is less than -4 and greater than or equal to 3

OpenStudy (anonymous):

thank youu! :)

OpenStudy (decentnabeel):

wellcome :) @_lexx.